OpenStudy (mary.rojas):

Please walk me through this : Given 2x-y=4 and (2^x)(2^y)=32, which of the following represents x-y : a. -2 b. -1 c. 0 d. 1 e. 2

OpenStudy (mary.rojas):

@jdoe0001 do you know how to do this?

OpenStudy (jdoe0001):

well keep in mind that \(\bf 2^x\times 2^y = 36\\ \color{blue}{36 = 2^5\\ a^x \times a^y = a^{x+y}}\\ \text{so the only 2 possible values lies within 5}\)

OpenStudy (mary.rojas):

you lost me at 36=2^5

OpenStudy (mary.rojas):

2^5 is 32

OpenStudy (jdoe0001):

then we'd just have to check that those 2 values give us 2x-y = 4 so let's try a combination 4 and 1 2(4)-(1) \(\bf \ne\) 4 another combination 3 and 2 \(\bf 2(3) -(2) = 4\) what about 2 and 3 2(2)-(3) \(\bf \ne\) 4

OpenStudy (jdoe0001):

so the one that worked was 3 and 2 so y = 2, x =3, and you'd know then what x-y is :)

OpenStudy (jdoe0001):

ohh .. \(\bf 2^5 = 36\) that is 2 x 2 x 2 x 2 x 2 = 36

OpenStudy (jdoe0001):

confused still?

OpenStudy (jdoe0001):

acck, I meant 32, not 36 :/

OpenStudy (jdoe0001):

so much for my typos :(

OpenStudy (mary.rojas):

ohhhh and 3,2 has to work for the other one right: (2^x)(2^y)=32 and i checked and it does....... oh ok at the 2^5 lol and the answer is -1?

OpenStudy (mary.rojas):

I MEAN +1 SORRY TYPO

OpenStudy (jdoe0001):

yes +1 :)

OpenStudy (jdoe0001):

\(\bf 2^x\times 2^y = 32\\ 2^x\times 2^y \implies 2^{x+y} = 2^5\\ x+y = 5\)

OpenStudy (mary.rojas):

thanks, that was actually easy :)

OpenStudy (jdoe0001):

yw